### v1

# On the error term in the explicit formula of Riemann–von Mangoldt II

Abstract.
We give an explicit $O(x/T)$ error term for the truncated Riemann–von Mangoldt explicit formula. For large $x$, this provides a modest improvement over previous work, which we demonstrate via an application to a result on primes between consecutive powers.
1. Introduction
This paper gives a new explicit estimate for the error in the truncated Riemann–von Mangoldt explicit formula. This is a sequel to [4], which gave a related result. In contrast to [4], this new estimate is asymptotically sharper, but allows for less control over the point of truncation.
Let $\psi(x)$ denote the Chebyshev prime-counting function with
$\psi(x)=\sum_{n\leq x}\Lambda(n)=\sum_{p^{k}\leq x}\log p,$
where $p^{k}$ are prime powers with any integer $k\geq 1$, and $\Lambda(n)$ is the von Mangoldt function. The truncated Riemann–von Mangoldt formula can be written as
$\psi(x)=x-\sum_{\begin{subarray}{c}\rho=\beta+i\gamma\\ |\gamma|\leq T\end{subarray}}\frac{x^{\rho}}{\rho}+E(x,T),$ (1.1)
where the sum is over all non-trivial zeros $\rho=\beta+i\gamma$ of the Riemann zeta-function $\zeta(s)$ that have $|\gamma|\leq T$, and $E(x,T)$ is an error term.
In [4, Thm 1.2], the authors prove the following.
Theorem 1.1.
For any $\alpha\in(0,1/2]$ there exist constants $M$ and $x_{M}$ such that for $\max\{51,\log x\}<T<(x^{\alpha}-2)/2$,
$\psi(x)=x-\sum_{\begin{subarray}{c}|\gamma|\leq T\end{subarray}}\frac{x^{\rho}}{\rho}+O^{*}\left(M\frac{x\log x}{T}\right)$ (1.2)
for all $x\geq x_{M}$. Some admissible values of $x_{M}$, $\alpha$ and $M$ are $(40,1/2,5.03)$ and $(10^{3},1/100,0.5597)$, with more given in [4, Table 4].
This result is reached via the Perron formula, which expresses the partial sums of an arithmetic function as an integral over its corresponding Dirichlet series. In particular, the truncated Perron formula can be written as (e.g., see [12, Thm. 5.2, Cor. 5.3] or [15, Chp. 4.4])
$\sum_{n\leq x}a_{n}=\frac{1}{2\pi i}\int_{\kappa-iT}^{\kappa+iT}F(s)\frac{x^{s}}{s}\mathrm{d}s+R(x,T).$ (1.3)
Here, $R(x,T)$ is an error term, and the formula is valid for $x>0$, any Dirichlet series $F(s)=\sum_{n=1}^{\infty}a_{n}n^{-s}$ that converges in $\Re(s)>\kappa_{a}$, any $\kappa>\max(\kappa_{a},0)$, and bounded $T>1$. The admissible range for the point of truncation $T$ is usually a function of $x$, and can be somewhat adjusted for explicit estimates.
In Section 3 of [19], Wolke gave a method to estimate the integral in (1.3). This was made explicit in [4, Thm. 4] for the case $F(s)=\sum_{n\geq 1}\Lambda(n)n^{-s}=-(\zeta^{\prime}/\zeta)(s)$ and $\kappa=1+1/\log x$. The authors then adapted a method of Ramaré in [17, Thm. 1.1, Thm. 2.1] to estimate $R(x,T)$, [4, Thm. 2.1]. Combining these two estimates in (1.3) resulted in Theorem 1.1.
In [17, Thm. 1.2, Thm. 3.1], Ramaré then proved a weighted and averaged version of (1.3) with an asymptotically sharper estimate for the error term. The purpose of this paper is to make this method of Ramaré explicit, and apply it in the context of the Riemann–von Mangoldt formula. Although [17, Thm. 3.1] is already explicit, there are several errors in the proofs and statements of Theorems 1.2 and 3.1 of [17] that we need to correct or circumvent. For instance, there is a missing factor of $k+1$ in the error term of [17, Thm. 3.1], a typo in the definition (3.1) of [17], and some measure-theoretic problems which were confirmed by Ramaré through private correspondence (see also [18]). After mending these errors, and simplifying the result for our purposes, our main theorem is as follows.
Theorem 1.2.
For any $\alpha\in(0,1/2]$ and $\omega\in[0,1]$ there exist constants $M$ and $x_{M}$ such that for some $T^{*}\in[T,2T]$ with $\max\{51,\log^{2}x\}<T<(x^{\alpha}-2)/4$, we have
$\psi(x)=x-\sum_{\begin{subarray}{c}|\gamma|\leq T^{*}\end{subarray}}\frac{x^{\rho}}{\rho}+O^{*}\left(M\frac{x(\log x)^{1-\omega}}{T}\right)$ (1.4)
for all $x\geq x_{M}$. Some admissible values of $x_{M}$, $\alpha$ and $M$ are given in Table 1.
Table 1. Some corresponding values of $x_{M}$, $\alpha$, $\lambda$ and $M$ for Theorem 1.2. Here $\lambda$ is a parameter which we optimise over in the proof of the theorem. The last two entries are specifically used to prove Theorem 1.4
$\log(x_{M})$ $\alpha$ $\omega$ $\lambda$ $M$
$40$ $1/2$ 00 $0.43$ $2.914$
$10^{3}$ $1/2$ 00 $0.42$ $1.706$
$10^{10}$ $1/2$ $0.3$ $0.42$ $2.275$
$10^{13}$ $1/2$ $1$ $10^{-4}$ $19.81$
$10^{3}$ $1/10$ $0.2$ $0.07$ $1.260$
$10^{10}$ $1/10$ $0.9$ $10^{-3}$ $4.431$
$10^{3}$ $1/100$ $0.8$ $0.07$ $3.615$
$10^{10}$ $1/100$ $1$ $10^{-3}$ $9.631$
$10^{3}$ $1/85$ $0.9$ $0.07$ $6.391$
$4\cdot 10^{3}$ $1/85$ $0.9$ $0.05$ $5.462$
Although we now have less control over the point of truncation, for most applications Theorem 1.2 implies that the factor of $\log x$ in (1.2) can be reduced or removed. At first glance, it may seem preferable to just take $\omega=1$. However, the corresponding $M$ may be such that the overall error estimate is larger for a wide range of $x$, compared to the $M$ one obtains with a smaller $\omega$. Hence, for small to moderate values of $x$ the best estimate will often come from using some $\omega<1$.
It would also be possible to sharpen the explicit result: we see two immediate avenues. In the proof of Theorem 1.2, one could set $\kappa$ to be another function such as $\kappa=1+c/\log x$ and optimise over $c$. This would also require reworking [4, Thm. 4] for such a choice of $\kappa$. Another potential improvement is to modify the range for $T^{*}$ (currently $[T,2T]$). This would amount to optimising the choice of the parameter $\xi>1$ which we introduce in §2. Either of these improvements would be rather systematic to pursue, but we have chosen not to as they would complicate and lengthen this paper for only a small to moderate improvement. In particular, neither of these improvements would improve the asymptotic form of (1.4). Moreover, the main error terms in the proof of Theorem 1.2 rely more on other results, such as the size of the zero-free region of $\zeta(s)$.
We also have the following variant of Theorem 1.2 for primes in intervals.
Theorem 1.3.
With the same notation as in Theorem 1.2, we have for all $x\geq x_{M}$ and any nonnegative function $h=h(x)$,
$\psi(x+h)-\psi(x)=h-$ $\sum_{|\gamma|\leq T^{*}}\frac{(x+h)^{\rho}-x^{\rho}}{\rho}$
$\quad+O^{*}\left(M\frac{x+h}{T}(\log(x+h))^{1-\omega}+M\frac{x}{T}(\log x)^{1-\omega}\right).$
To demonstrate the power of our new estimates we use Theorem 1.3 to obtain the following result.
Theorem 1.4.
There is at least one prime between $n^{90}$ and $(n+1)^{90}$ for all $n\geq 1$.
This improves upon the previous result in [4, Thm. 1.3], which found primes between consecutive $140^{\text{th}}$ powers.
1.1. Outline of paper
In Section 2 we make explicit a weighted averaged version of the truncated Perron formula. From this, we deduce a corollary in Section 3 for a “modified” version of the truncated Perron formula for general arithmetic functions. This corollary is then used in conjunction with [4, Thm. 4.1] to reach Theorem 1.2 in Section 4.
2. An averaged, weighted, truncated Perron formula
In this section we prove Theorem 2.1 stated below, which is an error estimate for a weighted averaged version of the truncated Perron formula. This result is a modification of [17, Thm. 3.1] and its proof also fixes several minor errors appearing in [17].
First, some notation: let $k,\xi\geq 1$. We call a function $w$, defined over $[1,\xi]$, $(k,\xi)$-admissible for $k\geq 1$ if it has the following properties:
  • (1)
    it is $k$-times differentiable, and $w^{(k)}$ is in $L^{1}$,
  • (2)
    $\int_{1}^{\xi}w(t)\mathrm{d}t=1$,
  • (3)
    $w^{(\ell)}(1)=w^{(\ell)}(\xi)=0$ for $20\leq\ell\leq k-$ 2. This condition is empty when $k=1$.
For such a function $w$, we define
$2\pi N_{k,\xi}(w)$ $=\frac{1}{\xi}|w^{(k-1)}(\xi)|+|w^{(k-1)}(1)|+k!\sum_{0\leq h\leq k}\frac{1}{h!}\int_{1}^{\xi}\frac{|w^{(h)}(u)|}{u}\mathrm{d}u,$
$\mathscr{L}_{\xi}(w)$ $=\int_{1}^{\xi}u|w(u)|\mathrm{d}u/\pi$
and $\theta^{\prime}=\theta^{\prime}_{k,\xi}(w)$ is the positive point of intersection of the two functions
$f_{1}(y)=\frac{2N_{k,\xi}(w)}{|y|^{k+1}}\quad\text{and}\quad f_{2}(y)=1+|y|L_{\xi}(w).$
We aim to prove the following theorem.
Theorem 2.1.
Let $k\geq 1$ be an integer and let $\xi>1$ be a real number. Let $w$ be a $(k,\xi)$-admissible function. Let $F(s)=\sum_{n\geq 1}a_{n}/n^{s}$ be a Dirichlet series that converges absolutely for $\Re(s)>\kappa_{a}$. For $x\geq 1$, $T\geq 1$, and $\kappa>\kappa_{a}>0$ we have
$\sum_{n\leq x}a_{n}$ $=\frac{1}{2\pi i}\int_{T}^{\xi T}\int_{\kappa-it}^{\kappa+it}F(s)\frac{x^{s}}{s}\mathrm{d}s\frac{w(t/T)}{T}\mathrm{d}t$
$+O^{*}\left((k+1)\frac{2N_{k,\xi}(w)}{T^{k+1}}\int_{\theta^{\prime}/T}^{\infty}\sum_{|\log(x/n)|\leq u}\frac{|a_{n}|}{n^{\kappa}}\frac{x^{\kappa}\mathrm{d}u}{u^{k+2}}\right).$ (2.1)
To prove Theorem 2.1 we need the following lemma that uses the function
$v(y)=\begin{cases}1\quad y\geq 0\\ 0\quad y<0.\end{cases}$
Lemma 2.2.
Let $w$ be a $(k,\xi)$-admissible function. For $\kappa^{\prime}>0$ and $y\in\mathbb{R}$,
$\left|v(y)-\frac{1}{2\pi i}\int_{1}^{\xi}\int_{\kappa^{\prime}-i\tau}^{\kappa^{\prime}+i\tau}\frac{e^{sy}}{s}\mathrm{d}sw(\tau)\mathrm{d}\tau\right|\leq\Delta(y),$
where
$\Delta(y)=\begin{cases}\min\left\{\frac{2e^{y\kappa^{\prime}}}{|y|^{k+1}}N_{k,\xi}(w),\left|v(y)-\frac{e^{y\kappa^{\prime}}}{\pi}\arctan(1/\kappa^{\prime})\right|+|y|e^{y\kappa^{\prime}}L_{\xi}(w)\right\},&y\neq 0,\\ \left|v(y)-\frac{e^{y\kappa^{\prime}}}{\pi}\arctan(1/\kappa^{\prime})\right|+|y|e^{y\kappa^{\prime}}L_{\xi}(w),&y=0.\end{cases}$ (2.2)
Proof.
We begin by noting that
$\int_{\kappa^{\prime}-i\tau}^{\kappa^{\prime}+i\tau}\frac{e^{ys}}{s}\mathrm{d}s=e^{y\kappa^{\prime}}\int_{\kappa^{\prime}-i\tau}^{\kappa^{\prime}+i\tau}\frac{1}{s}\mathrm{d}s+ie^{y\kappa^{\prime}}\int_{-\tau}^{\tau}\frac{e^{iyt}-1}{\kappa^{\prime}+it}\mathrm{d}t.$
The first integral is equal to $2i\text{arctan}(1/\kappa^{\prime})$. For $y\neq 0$, the second integral can be bounded by $2\tau|y|$ using the identity
$\left|\frac{e^{iyt}-1}{iyt}\right|=\left|\int_{0}^{1}e^{iytu}\mathrm{d}u\right|\leq 1,$
and note that the bound $|e^{iyt}-1|\leq|yt|$ can be separately verified for $y=0$. This leads to
$\left|v(y)-\frac{1}{2\pi i}\int_{1}^{\xi}\int_{\kappa^{\prime}-i\tau}^{\kappa^{\prime}+i\tau}\frac{e^{sy}}{s}\mathrm{d}sw(\tau)\mathrm{d}\tau\right|$ $\leq\left|v(y)-\frac{e^{y\kappa^{\prime}}}{\pi}\text{arctan}(1/\kappa^{\prime})\right|+|y|e^{y\kappa^{\prime}}L_{\xi}(w).$
for all $y\in\operatorname{\mathbb{R}}$.
To prove the other part of the $y\neq 0$ case in (2.2), we first consider $y<0$. For any $K>\kappa^{\prime}$ we have
$\left(\int_{\kappa^{\prime}-i\tau}^{\kappa^{\prime}+i\tau}+\int_{\kappa^{\prime}+i\tau}^{K+i\tau}+\int_{K+i\tau}^{K-i\tau}+\int_{K-i\tau}^{\kappa^{\prime}-i\tau}\right)\frac{e^{ys}}{s}\mathrm{d}s=0.$ (2.3)
We multiply both sides of (2.3) by $w(\tau)$ and integrate over $\tau$ from 1 to $\xi$. The third integral can be bounded by
$\left|\int_{1}^{\xi}\int_{K+i\tau}^{K-i\tau}\frac{e^{ys}}{s}\mathrm{d}sw(\tau)\mathrm{d}\tau\right|$ $=\int_{1}^{\xi}\int_{\tau}^{-\tau}\frac{ie^{y(K+it)}}{K+it}\mathrm{d}tw(\tau)\mathrm{d}\tau$
$\leq\int_{1}^{\xi}\int_{-\tau}^{\tau}\frac{e^{yK}}{K}\mathrm{d}t\left|w(\tau)\right|\mathrm{d}\tau\leq\max_{[1,\xi]}|w(\tau)|\cdot\frac{\xi^{2}e^{yK}}{K},$ (2.4)
which goes to zero as $K\to\infty$, noting that $\max_{[1,\xi]}|w(\tau)|$ exists because $w(\tau)$ is differentiable and thus continuous on $[1,\xi]$. We now bound the integral from $\kappa^{\prime}+i\tau$ to $K+i\tau$ in the limit $K\to\infty$, and note that the following argument also gives the same bound for the integral from $K-i\tau$ to $\kappa^{\prime}-i\tau$. We start by writing
$\int_{1}^{\xi}\int_{\kappa^{\prime}+i\tau}^{K+i\tau}\frac{e^{sy}}{s}\mathrm{d}sw(\tau)\mathrm{d}\tau=\int_{\kappa^{\prime}}^{K}e^{uy}\int_{1}^{\xi}\frac{e^{i\tau y}w(\tau)}{u+i\tau}\mathrm{d}\tau\mathrm{d}u.$ (2.5)
To deal with the inner integral, we set $f(\tau)=w(\tau)/(u+i\tau)$ and apply the Leibnitz formula for the derivative of a product. In particular, for any $m\geq 0$ we have
$f^{(m)}(\tau)=\sum_{0\leq h\leq m}\binom{m}{h}\frac{i^{m-h}(m-h)!w^{(h)}(\tau)}{(u+i\tau)^{m-h+1}}.$
Since $w^{(\ell)}(1)=w^{(\ell)}(\xi)=0$ for $20\leq\ell\leq k-$ 2, this implies
$f^{(k-1)}(1)=\frac{w^{(k-1)}(1)}{u+i},\qquad f^{(k-1)}(\xi)=\frac{w^{(k-1)}(\xi)}{u+i\xi}.$
Integrating by parts $k$ times thus gives
$\int_{1}^{\xi}\frac{e^{i\tau y}w(\tau)}{u+i\tau}\mathrm{d}\tau$ $=\frac{(-1)^{(k-1)}f^{(k-1)}(\xi)e^{\xi iy}}{(iy)^{k}}-\frac{(-1)^{k-1}f^{(k-1)}(1)e^{iy}}{(iy)^{k}}$
$\qquad+\frac{(-1)^{k}}{(iy)^{k}}\int_{1}^{\xi}e^{i\tau y}f^{(k)}(\tau)\mathrm{d}\tau.$
Substituting this into (2.5) gives
$\int_{1}^{\xi}\int_{\kappa^{\prime}+i\tau}^{K+i\tau}\frac{e^{sy}}{s}\mathrm{d}sw(\tau)\mathrm{d}\tau=$
$\qquad\frac{(-1)^{k-1}w^{(k-1)}(\xi)}{(iy)^{k}}\int_{\kappa^{\prime}}^{K}\frac{e^{(u+\xi i)y}}{u+i\xi}\mathrm{d}u-\frac{(-1)^{k-1}w^{(k-1)}(1)}{(iy)^{k}}\int_{\kappa^{\prime}}^{K}\frac{e^{(u+i)y}}{u+i}\mathrm{d}u$
$\qquad+\sum_{0\leq h\leq k}\binom{k}{h}\int_{1}^{\xi}\frac{(-1)^{k}i^{k-h}(k-h)!w^{(h)}(\tau)}{(iy)^{k}}\int_{\kappa^{\prime}}^{K}\frac{e^{(u+i\tau)y}}{(u+i\tau)^{k-h+1}}\mathrm{d}u\mathrm{d}\tau.$ (2.6)
For the first integral, we bound $1/(u+i\xi)$ by $1/\xi$ and integrate $e^{uy}$ to obtain
$\left|\frac{(-1)^{k-1}w^{(k-1)}(\xi)}{(iy)^{k}}\int_{\kappa^{\prime}}^{K}\frac{e^{(u+\xi i)y}}{u+i\xi}\mathrm{d}u\right|\leq\frac{|w^{(k-1)}(\xi)|}{|y|^{k}}\cdot\frac{e^{\kappa^{\prime}y}}{\xi|y|}$
since $K\to\infty$. The other integrals in (2.6) are bounded in analogous ways to give
$\int_{1}^{\xi}\int_{\kappa^{\prime}+i\tau}^{K+i\tau}\frac{e^{sy}}{s}\mathrm{d}sw(\tau)\mathrm{d}\tau\leq\frac{2\pi e^{\kappa^{\prime}y}N_{k,\xi}(w)}{|y|^{k+1}}.$
Using this bound in (2.3) gives
$\left|v(y)-\frac{1}{2\pi i}\int_{1}^{\xi}\int_{\kappa^{\prime}-i\tau}^{\kappa^{\prime}+i\tau}\frac{e^{ys}}{s}\mathrm{d}sw(\tau)\mathrm{d}\tau\right|\leq\frac{2e^{y\kappa^{\prime}}}{|y|^{k+1}}N_{k,\xi}(w).$
The case $y>0$ is almost identical, but for $K<0$ we begin with
$\left(\int_{\kappa^{\prime}-i\tau}^{\kappa^{\prime}+i\tau}+\int_{\kappa^{\prime}+i\tau}^{K+i\tau}+\int_{K+i\tau}^{K-i\tau}+\int_{K-i\tau}^{\kappa^{\prime}-i\tau}\right)\frac{e^{ys}}{s}\mathrm{d}s=2\pi i.$ (2.7)
After multiplying both sides of (2.7) by $w(\tau)$ and integrating over $\tau$ from 1 to $\xi$, we find that the third integral approaches 0 as $K\to-\infty$, as it has the same bound as in (2). For the integrals from $\kappa^{\prime}+i\tau$ to $K+i\tau$ and from $K-i\tau$ to $\kappa^{\prime}-i\tau$, we have the same bounds as before in the limit $K\to-\infty$. Using these estimates in (2.7) gives
$\left|2\pi i-\int_{1}^{\xi}\int_{\kappa^{\prime}-i\tau}^{\kappa^{\prime}+i\tau}\frac{e^{ys}}{s}\mathrm{d}sw(\tau)\mathrm{d}\tau\right|\leq\frac{4\pi e^{y\kappa^{\prime}}}{y^{k+1}}N_{k,\xi}(w),$
and thus
$\left|v(y)-\frac{1}{2\pi i}\int_{1}^{\xi}\int_{\kappa^{\prime}-i\tau}^{\kappa^{\prime}+i\tau}\frac{e^{ys}}{s}\mathrm{d}sw(\tau)\mathrm{d}\tau\right|\leq\frac{2e^{y\kappa^{\prime}}}{y^{k+1}}N_{k,\xi}(w).\qed$
Proof of Theorem 2.1.
We now prove Theorem 2.1. To do this, we need to bound
$\left|\sum_{n\geq 1}a_{n}v(T\log(x/n))-\frac{1}{2\pi iT}\int_{T}^{\xi T}\int_{\kappa-it}^{\kappa+it}F(s)\frac{x^{s}\mathrm{d}s}{s}w(t/T)\mathrm{d}t\right|x^{-\kappa}$
for any $x\geq 1$, $T>0$, and $\kappa>\kappa_{a}>0$, where $\kappa_{a}$ is the abscissa of absolute convergence of $F(s)$. First, we take $\kappa=\kappa^{\prime}T$ for any $\kappa^{\prime}>0$, and use the substitutions $s=\tilde{s}T$ and $t=\tau T$ to obtain
$\left|\sum_{n\geq 1}a_{n}v(T\log(x/n))-\frac{1}{2\pi iT}\int_{T}^{\xi T}\int_{\kappa-it}^{\kappa+it}F(s)\frac{x^{s}\mathrm{d}s}{s}w(t/T)\mathrm{d}t\right|x^{-\kappa}$
$\qquad\leq\sum_{n\geq 1}\frac{|a_{n}|}{n^{\kappa}}\left|v(T\log(x/n))-\frac{1}{2\pi iT}\int_{T}^{\xi T}\int_{\kappa-it}^{\kappa+it}\left(\frac{x}{n}\right)^{s}\frac{\mathrm{d}s}{s}w(t/T)\mathrm{d}t\right|\left(\frac{n}{x}\right)^{\kappa}$
$\qquad=\sum_{n\geq 1}\frac{|a_{n}|}{n^{\kappa}}\left|v(T\log(x/n))-\frac{1}{2\pi iT}\int_{1}^{\xi}\int_{\kappa^{\prime}-i\tau}^{\kappa^{\prime}+i\tau}e^{T\log\left(\frac{x}{n}\right)\tilde{s}}\frac{\mathrm{d}\tilde{s}}{\tilde{s}}w(\tau)\mathrm{d}\tau\right|e^{-\kappa^{\prime}T\log(x/n)}.$ (2.8)
Next, we apply Lemma 2.2. For $\theta>0$ we let
$c_{\theta}=\max_{0<|y|<\theta}\Delta(y),$ (2.9)
where $\Delta(y)$ is as defined in (2.2). Splitting the sum at $\theta$ in (2) and taking $y=T\log(x/n)$ in Lemma 2.2 gives
$\left|\sum_{n\geq 1}a_{n}v(T\log(x/n))-\frac{1}{2\pi iT}\int_{T}^{\xi T}\int_{\kappa-it}^{\kappa+it}F(s)\frac{x^{s}\mathrm{d}s}{s}w(t/T)\mathrm{d}t\right|x^{-\kappa}$
$\qquad\leq c_{\theta}\sum_{T|\log(x/n)|<\theta}\frac{|a_{n}|}{n^{\kappa}}+\frac{2N_{k,\xi}(w)}{T^{k+1}}\sum_{T|\log(x/n)|\geq\theta}\frac{|a_{n}|}{n^{\kappa}|\log(x/n)|^{k+1}}.$
Now, since
$\sum_{T|\log(x/n)|\geq\theta}\frac{|a_{n}|}{n^{\kappa}|\log(x/n)|^{k+1}}$
$\qquad=\sum_{T|\log(x/n)|\geq\theta}\frac{|a_{n}|}{n^{\kappa}}\int_{|\log(x/n)|}^{\infty}\frac{(k+1)}{u^{k+2}}\mathrm{d}u$
$\qquad=(k+1)\int_{\theta/T}^{\infty}\sum_{\theta/T\leq|\log(x/n)|\leq u}\frac{|a_{n}|}{n^{\kappa}}\frac{\mathrm{d}u}{u^{k+2}}$
$\qquad=(k+1)\int_{\theta/T}^{\infty}\sum_{|\log(x/n)|\leq u}\frac{|a_{n}|}{n^{\kappa}}\frac{\mathrm{d}u}{u^{k+2}}-\frac{T^{k+1}}{\theta^{k+1}}\sum_{|\log(x/n)|<\theta/T}\frac{|a_{n}|}{n^{\kappa}}$
we have
$\left|\sum_{n\leq x}a_{n}v(T\log(x/n))-\frac{1}{2\pi iT}\int_{T}^{\xi T}\int_{\kappa-it}^{\kappa+it}F(s)\frac{x^{s}\mathrm{d}s}{s}w(t/T)\mathrm{d}t\right|x^{-\kappa}$
$\qquad\leq(k+1)\frac{2N_{k,\xi}(w)}{T^{k+1}}\int_{\theta/T}^{\infty}\sum_{|\log(x/n)|\leq u}\frac{|a_{n}|}{n^{\kappa}}\frac{\mathrm{d}u}{u^{k+2}}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\left(c_{\theta}-\frac{2N_{k,\xi}(w)}{\theta^{k+1}}\right)\sum_{|\log(x/n)|<\theta/T}\frac{|a_{n}|}{n^{\kappa}}.$ (2.10)
Note that $\left|v(y)e^{-y\kappa^{\prime}}-\frac{1}{\pi}\arctan(1/\kappa^{\prime})\right|<1$, so taking $\theta=\theta^{\prime}$ we have
$c_{\theta^{\prime}}=\frac{2N_{k,\xi}(w)}{(\theta^{\prime})^{k+1}}$
and the second term in (2) vanishes. This proves the theorem. ∎
3. A modified Perron formula
As Theorem 2.1 involves an integral from $T$ to $\xi T$, we can set $\xi=2$ and use a continuity argument to deduce a strong bound for some $T^{*}\in[T,2T]$. First, we need to choose a weight function $w(t)$. Roughly following [17, Sect. 5], we define
$f(v)=\begin{cases}v(1-v),&\text{when $v\in[0,1]$},\\ 0,&\text{else},\end{cases}$
and
$w(u)=6f\left(u-1\right).$
Note that
$\int_{1}^{2}w(u)\mathrm{d}u=1$ (3.1)
and $w(u)$ is a $(1,2)$-admissible function. The definitions in Section 2 also imply that for $k=1$ and $\xi=2$ we have
$\mathscr{L}_{2}(w)=\frac{3}{2\pi},$
$0.4460\leq N_{1,2}(w)\leq 0.4461$ (3.2)
and
$0.8029\leq\theta^{\prime}_{1,2}(w)\leq 0.8030$
We now substitute this choice of $w(u)$ into Theorem 2.1, giving the following result.
Corollary 3.1.
Let $w$ and $\theta^{\prime}=\theta^{\prime}_{1,2}(w)$ be as above. Let $x\geq 1$ and $T>1$. Let $F(s)=\sum_{n\geq 1}a_{n}n^{-s}$ be a Dirichlet series and $\kappa>0$ be a real parameter larger than $\kappa_{a}$ — the abscissa of absolute convergence of $F$. For any $\lambda\geq\theta^{\prime}/T$, there exists a $T^{*}\in[T,2T]$ such that
$\sum_{n\leq x}a_{n}=\frac{1}{2\pi i}\int_{\kappa-iT^{*}}^{\kappa+iT^{*}}F(s)\frac{x^{s}}{s}\mathrm{d}s+O^{*}(\mathfrak{R}),$ (3.3)
where $\mathfrak{R}=\mathfrak{R}(x,T,\kappa)$ satisfies
$\mathfrak{R}\leq\frac{1.785}{T^{2}}\left(\frac{x^{\kappa}}{2\lambda^{2}}\sum_{n\geq 1}\frac{|a_{n}|}{n^{\kappa}}+e^{\kappa\lambda}\int_{\theta^{\prime}/T}^{\lambda}\sum_{|\log(x/n)|\leq u}\frac{|a_{n}|\mathrm{d}u}{u^{3}}\right).$ (3.4)
Proof.
Let $\xi=2$ and $k=1$ in Theorem 2.1, and set
$\mathfrak{R}^{\prime}=\frac{4N_{1,2}(w)}{T^{2}}\int_{\theta^{\prime}/T}^{\infty}\sum_{|\log(x/n)|\leq u}\frac{|a_{n}|}{n^{\kappa}}\frac{x^{\kappa}\mathrm{d}u}{u^{3}}$ (3.5)
to be the error term in (2.1). A change of variables in (3.1) gives
$\int_{T}^{2T}\frac{w(t/T)}{T}\mathrm{d}t=1.$ (3.6)
Thus, (2.1) can be rewritten as
$\int_{T}^{2T}\left(\sum_{n\leq x}a_{n}-\frac{1}{2\pi i}\int_{\kappa-it}^{\kappa+it}F(s)\frac{x^{s}}{s}\mathrm{d}s\right)\frac{w(t/T)}{T}\mathrm{d}t=O^{*}(\mathfrak{R}^{\prime}).$ (3.7)
Let $\epsilon=1-1/2\pi$ and
$G(x,t)=\sum_{n\leq x}a_{n}-\frac{1}{2\pi i}\int_{\kappa-it}^{\kappa+it}F(s)\frac{x^{s}}{s}\mathrm{d}s.$ (3.8)
Given (3.6) and that $G(x,t)$ is continuous in $t$, (3.7) implies that there exists a $T^{*}\in[T,2T]$ satisfying
$\left|G(x,T^{*})\right|\leq\mathfrak{R}^{\prime}.$ (3.9)
Splitting the integral in (3.5) at $u=\lambda$ and using $N_{1,2}(w)\leq 0.4461$ gives
$\mathfrak{R}^{\prime}$ $\leq\frac{4\cdot 0.4461}{T^{2}}\left(\frac{x^{\kappa}}{2\lambda^{2}}\sum_{n\geq 1}\frac{|a_{n}|}{n^{\kappa}}+\int_{\theta^{\prime}/T}^{\lambda}\sum_{|\log(x/n)|\leq u}|a_{n}|\left(\frac{x}{n}\right)^{\kappa}\frac{1}{u^{3}}\mathrm{d}u\right)$
$\leq\frac{1.785}{T^{2}}\left(\frac{x^{\kappa}}{2\lambda^{2}}\sum_{n\geq 1}\frac{|a_{n}|}{n^{\kappa}}+e^{\kappa\lambda}\int_{\theta^{\prime}/T}^{\lambda}\sum_{|\log(x/n)|\leq u}\frac{|a_{n}|}{u^{3}}\mathrm{d}u\right),$ (3.10)
as desired. ∎
We also have the following variant of Corollary 3.1 for sums $\sum_{x<n\leq x+h}a_{n}$.
Corollary 3.2.
Keep the notation of Corollary 3.1. For any nonnegative function $h=h(x)$, we have
$\sum_{x<n\leq x+h}a_{n}=\frac{1}{2\pi i}\int_{\kappa-iT^{*}}^{\kappa+iT^{*}}F(s)\frac{(x+h)^{s}-x^{s}}{s}\mathrm{d}s+O^{*}(\mathfrak{R}(x+h,T,\kappa)+\mathfrak{R}(x,T,\kappa)).$ (3.11)
Proof.
Repeat the argument in the proof of Corollary 3.1 for
$\sum_{x<n\leq x+h}a_{n}=\sum_{n\leq x+h}a_{n}-\sum_{n\leq x}a_{n}.\qed$
Note that simply applying Corollary 3.1 twice is not sufficient to prove Corollary 3.2. This is because we require $T^{*}\in[T,2T]$ to be the same value in both the integrals
$\int_{\kappa-iT^{*}}^{\kappa+iT^{*}}F(s)\frac{(x+h)^{s}}{s}\mathrm{d}s\qquad\text{and}\qquad\int_{\kappa-iT^{*}}^{\kappa+iT^{*}}F(s)\frac{(-x^{s})}{s}\mathrm{d}s.$
4. Proof of theorems 1.2 and 1.3
To prove Theorem 1.2 we use Corollary 3.1 for $a_{n}=\Lambda(n)$ and $\kappa=1+1/\log x$. Namely, for $x\geq x_{M}\geq e^{40}$ and $\max\{51,\log^{2}x\}<T<\frac{x^{\alpha}-2}{4}$ with any $\alpha\in(0,1/2]$ we have that for some $T^{*}\in[T,2T]$
$\psi(x)=\frac{1}{2\pi i}\int_{\kappa-iT^{*}}^{\kappa+iT^{*}}\left(-\frac{\zeta^{\prime}}{\zeta}(s)\right)\frac{x^{s}}{s}\mathrm{d}s$
$\qquad\qquad\qquad+O^{*}\left(\frac{1.785}{T^{2}}\left(\frac{x^{\kappa}}{2\lambda^{2}}\sum_{n\geq 1}\frac{\Lambda(n)}{n^{\kappa}}+e^{\kappa\lambda}\int_{\theta^{\prime}/T}^{\lambda}\sum_{|\log(x/n)|\leq u}\frac{\Lambda(n)}{u^{3}}\right)\mathrm{d}u\right),$ (4.1)
where $\lambda$ is a free parameter and $\theta^{\prime}$ is as defined in Corollary 3.1.
For the first term, Theorem B.1 in Appendix B implies that for any $\omega\in[0,1]$ there exists a constant $K$ corresponding to $x_{M}$ (given as $x_{K}$ in Table 2) such that
$\frac{1}{2\pi i}\int_{\kappa-iT^{*}}^{\kappa+iT^{*}}\left(-\frac{\zeta^{\prime}}{\zeta}(s)\right)\frac{x^{s}}{s}\mathrm{d}s=x-\sum_{|\gamma|\leq T^{*}}\frac{x^{\rho}}{\rho}+O^{*}\left(K\frac{x}{T}(\log x)^{1-\omega}\right).$ (4.2)
Next, by the main theorem in [6], the first sum in (4) is bounded by
$1.785\cdot\frac{x^{\kappa}}{2\lambda^{2}T^{2}}\sum_{n\geq 1}\frac{\Lambda(n)}{n^{\kappa}}\leq 1.785\cdot\frac{ex\log x}{2\lambda^{2}T^{2}}.$ (4.3)
For the second term in (4), the condition $|\log(x/n)|\leq u$ is equivalent to $xe^{-u}\leq n\leq xe^{u}$. For all $u\in[0,\lambda]$ we have $e^{-u}\geq 1-u$ and $e^{u}\leq 1+(c_{0}-1)u$ with
$c_{0}=c_{0}(\lambda)=\frac{e^{\lambda}-1}{\lambda}+1.$
Hence
$\frac{1.785e^{\kappa\lambda}}{T^{2}}\int_{\theta^{\prime}/T}^{\lambda}\sum_{|\log(x/n)|\leq u}\frac{\Lambda(n)}{u^{3}}\mathrm{d}u$ $\leq\frac{1.785e^{\kappa\lambda}}{T^{2}}\int^{\lambda}_{\theta^{\prime}/T}\frac{1}{u^{3}}\sum_{I(x,u)}\Lambda(n)\mathrm{d}u,$ (4.4)
where
$I(x,u)=\left\{n\geq 1\>:\>x-ux\leq n\leq x+(c_{0}-1)ux\right\}.$
Let $x_{-}=\max\{x-ux,0\}$ and $x_{+}=x+(c_{0}-1)ux$ for $\theta^{\prime}/T\leq u\leq\lambda$. Defining
$\theta(x)=\sum_{p\leq x}\log p,$
we have by an explicit form of the Brun–Titchmarsh theorem [13, Thm. 2],
$\theta(x_{+})-\theta(x_{-})$ $\leq\frac{2\log x_{+}}{\log(x_{+}-x_{-})}(x_{+}-x_{-})$
$\leq\frac{2\log(x+(c_{0}-1)ux)}{\log(c_{0}ux)}c_{0}ux$
$\leq c_{0}ux\cdot\mathscr{E}_{1}(x,T),$
where
$\mathscr{E}_{1}(x,T)=\frac{2\log(x+(c_{0}-1)\lambda x)}{\log(c_{0}\theta^{\prime}x/T)}.$
To obtain the corresponding inequality for $\psi$, we use [2, Cor. 5.1], which states that for all $x\geq e^{40}$,
$0<\psi(x)-\theta(x)<a_{1}x^{1/2}+a_{2}x^{1/3},$
with $a_{1}=1+1.93378\cdot 10^{-8}$ and $a_{2}=1.0432$. Thus,
$\sum_{n\in I(x,u)}\Lambda(n)$ $\leq\psi(x_{+})-\psi(x_{-})+\log x$
$<c_{0}ux\cdot\mathscr{E}_{1}(x,T)+a_{1}\sqrt{x_{+}}+a_{2}\sqrt[3]{x_{+}}+\log x$
$\leq c_{0}ux\cdot\mathscr{E}_{1}(x,T)+\mathscr{E}_{2}(x),$
where
$\mathscr{E}_{2}(x)=a_{1}(x+(c_{0}-1)\lambda x)^{1/2}+a_{2}(x+(c_{0}-1)\lambda x)^{1/3}+\log x.$
Substituting this into (4.4), we obtain
$\frac{1.785e^{\kappa\lambda}}{T^{2}}\int^{\lambda}_{\theta^{\prime}/T}\frac{1}{u^{3}}\sum_{I(x,u)}\Lambda(n)\mathrm{d}u$ $\leq\frac{1.785e^{\kappa\lambda}}{T^{2}}\int_{\theta^{\prime}/T}^{\lambda}\left(\frac{c_{0}x\cdot\mathscr{E}_{1}(x,T)}{u^{2}}+\frac{\mathscr{E}_{2}(x)}{u^{3}}\right)\mathrm{d}u$
$\leq 1.785e^{\kappa\lambda}\left(\frac{c_{0}x\cdot\mathscr{E}_{1}(x,T)}{\theta^{\prime}T}+\frac{\mathscr{E}_{2}(x)}{2(\theta^{\prime})^{2}}\right).$ (4.5)
Hence, for all $x\geq x_{M}$ and some $T^{*}\in[T,2T]$ we have
$\psi(x)=x-\sum_{|\gamma|\leq T^{*}}\frac{x^{\rho}}{\rho}+O^{*}\left(K\frac{x}{T}(\log x)^{1-\omega}+1.785\cdot\frac{ex\log x}{2\lambda^{2}T^{2}}\right)$
$\qquad\qquad\qquad+O^{*}\left(1.785e^{\kappa\lambda}\left(\frac{c_{0}x\cdot\mathscr{E}_{1}(x,T)}{\theta^{\prime}T}+\frac{\mathscr{E}_{2}(x)}{2(\theta^{\prime})^{2}}\right)\right).$ (4.6)
Finally, to obtain Theorem 1.2 and the values in Table 1, we optimise over $\lambda$ for specific $x_{M}$ and $\omega$, using $K$ from [4, Table 3].
Theorem 1.3 is proven in the same manner, using Corollary 3.2 in place of Corollary 3.1.
5. Application: primes between consecutive powers
In this section we prove Theorem 1.4, which is an explicit result on primes between consecutive powers that improves [3, Thm. 1] and [4, Thm. 1.3]. We use a similar method to that in [3] and [4], but use the new results for $\psi(x)$ in Theorem 1.2 and Theorem 1.3. We also make use of some recent improvements to zero-free regions for the Riemann zeta-function (listed in Appendix A).
To find primes in the interval $(n^{m},(n+1)^{m})$, we redefine $n=x^{\frac{1}{m}}$ and consider the slightly smaller interval $(x,x+h]$ with $h=mx^{1-1/m}$. Let $x_{M}$, $\alpha$, $\omega$, $\lambda$ and $M$ be as in Theorem 1.2. By Theorem 1.3, for $x\geq x_{M}$ and $\max\{51,\log^{2}x\}<T<(x^{\alpha}-2)/4$, there exists a $T^{*}\in[T,2T]$ for which
$\psi(x+h)-\psi(x)\geq\ h-\left|\sum_{|\gamma|\leq T^{*}}\frac{(x+h)^{\rho}-x^{\rho}}{\rho}\right|-M\frac{G(x,h)}{T}$ (5.1)
where $G(x,h)=(x+h)(\log(x+h))^{1-\omega}+x(\log x)^{1-\omega}$. If the right-hand side of (5.1) is positive for some range of $x$, there must be primes in $(x,x+h]$ for those $x$.
To estimate (5.1), we first bound the sum over zeta zeros. The following bounds are given in the proof of [3, Thm. 1], so some intermediate steps are omitted. We have
$\left|\frac{(x+h)^{\rho}-x^{\rho}}{\rho}\right|\leq hx^{\beta-1},$
where $\beta\in(0,1)$ is the real part of $\rho$. Hence,
$\left|\sum_{|\gamma|\leq T^{*}}\frac{(x+h)^{\rho}-x^{\rho}}{\rho}\right|\leq h\sum_{|\gamma|\leq 2T}x^{\beta-1}.$
It is then possible to write
$\sum_{|\gamma|\leq 2T}x^{\beta-1}=2\int_{0}^{1}x^{\sigma-1}\log x\sum_{\begin{subarray}{c}0<\gamma<2T\\ \beta>\sigma\end{subarray}}1\ d\sigma+2x^{-1}\sum_{0<\gamma<2T}1,$ (5.2)
wherein we can use bounds for $N(T)$ and $N(\sigma,T)$, which respectively denote the zero-counting function and zero-density function of $\zeta(s)$. We can also use zero-free regions for $\zeta(s)$ to curtail the upper end-point of the integral.
The most recent estimate for $N(T)$ is given by [8, Thm. 1.1,(1.4)]; for $N(\sigma,T)$ the most recent is in [11, Thm. 1.1,(1.7)], with updated constants in [10, Table 3]. For the zero-free region, we set $\nu(T)$ to be a combination of results: see (A.2) of Appendix A.
The same working as in the proof of [3, Thm. 1] follows, except we choose $T=x^{\mu}/2$ for any $\mu\in(0,1)$ such that $T$ satisfies $\max\{51,\log^{2}x\}\leq T\leq(x^{\alpha}-2)/4$. The factor of $1/2$ in the choice of $T$ is to account for the sum in (5.2) being over $|\gamma|\leq 2T$. The reasoning concludes with the condition that there are primes in $(x,x+h]$ for all $x\geq x_{M}$ satisfying (c.f. [3, Eq. (20)])
$1-F(x)-2M\frac{G(x,h)}{x^{\mu}h}+\frac{E(x)}{h}>0,$ (5.3)
where $F(x)$ is defined in [3, Eq. (19)] (but here we use $\mu$ in place of $\alpha$) and $E(x)$ is as in [3, Eq. (20)]. Note that our zero-free region function $\nu(T)$ (see (A.2)) appears in the definition of $F(x)$, and updates the analogous function in [3].
The goal is to find the smallest $m$ for which (5.3) holds for all $x\geq x_{0}$ and sufficiently small $x_{0}$. This involves optimising over $\mu$, and is similar to the process in [3, Sect. 4,pg. 114]. For any $\mu>1/m$, the left-hand side of (5.3) increases to 1 for sufficiently large $x$. The largest-order term is in $F(x)$, and is determined by the zero-free region. If (5.3) is true at the point where the zero-free region switches from using Lemma A.3 to Lemma A.4, then (5.3) will hold for all $x$ beyond this. There may also be a range of $\mu$ for which this is true, so the best choice of $\mu$ will keep (5.3) positive for the widest range of $x$ below this zero-free region crossover point.
We follow the previous steps for $m=90$, as this appears to be the smallest $m$ we can achieve with the given constants. From Table 1 we can use $M=6.555$ over $\log x\geq 10^{3}$, and $M=5.602$ over $\log x\geq 4\cdot 10^{3}$, each with $\omega=0.9$. Numerically optimising over $\mu$, we find that for $\mu=0.0113$, (5.3) is true for $10^{3}\leq x\leq 4\cdot 10^{3}$, and with $\mu=0.0112$ (5.3) holds for $x\geq 4\cdot 10^{3}$.
There are a few things to note about this result. First, as was done in [3], we used $\sigma_{1}=0.6$ in $F(x)$, as the parameter has a negligible effect on the results. Second, the values of $M$ use $\alpha=1/85$, which restricts $T<(x^{1/85}-2)/4$ by Theorem 1.2. As we take $T=x^{\mu}/2$, we can check that $\mu=0.0113$ implies Theorem 1.2 holds for $\log x\geq 1492$. Fortunately, the long-interval estimates for primes in [5] for $x\geq 4\cdot 10^{18}$ and $x\geq e^{600}$ verify that there are primes between consecutive $90^{\text{th}}$ powers for $\log(4\cdot 10^{18})\leq\log x\leq 2767$. The calculations of [Oliveira] verify this consecutive powers interval for the remaining $x\geq 1$. This proves Theorem 1.4.
Acknowledgements
Thanks to our supervisor, Tim Trudgian for his ideas and support when writing this paper. We also thank Olivier Ramaré for his correspondence regarding his earlier paper.
Appendix A Current zero-free regions
We use a zero-free region composed of the following four results.
Lemma A.1.
[14, Thm. 1.3] For $|t|\geq 2$ there are no zeros of $\zeta(\beta+it)$ in the region $\beta\geq 1-\nu_{1}(t)$, where
$\nu_{1}(t)=\frac{1}{5.558691\log|t|}.$
Lemma A.2.
[7, Thm. 3] For $|t|\geq 3$ there are no zeros of $\zeta(\beta+it)$ in the region $\beta\geq 1-\nu_{2}(t)$, where
$\nu_{2}(t)=\frac{1}{R(|t|)\log|t|},$
with
$R(t)=\frac{J(t)+0.685+0.155\log\log t}{\log t\left(0.04962-\frac{0.0196}{J(t)+1.15}\right)},$
and
$J(t)=\frac{1}{6}\log t+\log\log t+\log(0.618).$
Remark.
The expression for $J(t)$ in Lemma A.2 has been updated from [7] by using a more recent bound [9, Theorem 1.1].
Lemma A.3.
[20, Cor. 1.2] For $|t|\geq 3$ there are no zeros of $\zeta(\beta+it)$ in the region $\beta\geq 1-\nu_{3}(t)$, where
$\nu_{3}(t)=\frac{\log\log|t|}{21.233\log|t|}.$
Lemma A.4.
[1, Thm. 1.2] For $|t|\geq 3$ there are no zeros of $\zeta(\beta+it)$ in the region $\beta\geq 1-\nu_{4}(t)$ whereaaaThe constant 53.989 is actually better than what appears in the arXiv version of [1]. However, [1] has been recently updated by the author and a new version will appear shortly.
$\nu_{4}(t)=\frac{1}{53.989\log^{2/3}|t|(\log\log|t|)^{1/3}}.$ (A.1)
To get the widest zero-free region at any given $T$ we set
$\nu(T)=\begin{cases}\frac{1}{2},&\text{if $|T|\leq 3\cdot 10^{12}$},\\ \max\{\nu_{1}(T),\nu_{2}(T),\nu_{3}(T),\nu_{4}(T)\}&\text{otherwise,}\end{cases}$ (A.2)
where the constant $3\cdot 10^{12}$ comes from Platt and Trudgian’s computational verification of the Riemann hypothesis [16]. Note that $\nu_{2}(t)\geq\nu_{1}(t)$ for $t\geq\exp(46.3)$, $\nu_{3}(t)\geq\nu_{2}(t)$ for $t\geq\exp(170.3)$ and $\nu_{4}(t)\geq\nu_{3}(t)$ for $t\geq\exp(482036)$.
Appendix B Updating bounds for an integral over $\zeta^{\prime}/\zeta$
In [4, Section 4] the following theorem is proven.
Theorem B.1 ([4, Theorem 4.1]).
Let $\alpha\in(0,1]$ and $\omega\in[0,1]$. There exists constants $K$ and $x_{K}$ such that if $x\geq x_{K}$ and $\max\{51,\log x\}<T<\frac{x^{\alpha}}{2}-1$,
$\frac{1}{2\pi i}\int_{1+\varepsilon-iT}^{1+\varepsilon+iT}\left(-\frac{\zeta^{\prime}}{\zeta}(s)\right)\frac{x^{s}}{s}\mathrm{d}s=x-\sum_{\begin{subarray}{c}\rho=\beta+i\gamma\\ |\gamma|\leq T\end{subarray}}\frac{x^{\rho}}{\rho}+O^{*}\left(\frac{Kx}{T}(\log x)^{1-\omega}\right),$ (B.1)
where $\varepsilon=1/\log x$.
Here, corresponding values of $x_{K}$, $\alpha$, $\omega$ and $K$ are given in [4, Table 3]. A key component in computing values of $K$ is the use of explicit zero-free regions. As there have recently been improvements to explicit zero-free regions (Appendix A), the values in [4, Table 3] can be updated. We define $\nu(T)$ to be as in (A.2) and then use this updated zero-free region in the bounds (4.19) and (4.20) in [4]. This gives us the updated table below (Table 2), with values for $x_{K}$, $\alpha$, $\omega$ and $K$ that can be used in Theorem B.1.
Table 2. Some corresponding values of $x_{K}$, $\alpha$, $\omega$ and $K$ for Theorem B.1 that updates [4, Table 3]. Note that $\overline{\omega}$ and $D$ are parameters one optimises over in the original calculation in [4].
$\log(x_{K})$ $\alpha$ $\omega$ $\overline{\omega}$ $D$ $K$
$40$ $1/2$ 00 00 $2.053$
$10^{3}$ $1/2$ 00 00 $1.673$
$10^{10}$ $1/2$ $0.3$ $0.3$ $0.54$ $2.274$
$10^{13}$ $1/2$ $1$ $1.4$ $0.50$ $0.6367$
$10^{3}$ $1/10$ $0.2$ $0.2$ $0.35$ $1.215$
$10^{10}$ $1/10$ $0.9$ $0.9$ $0.53$ $3.441$
$10^{3}$ $1/100$ $0.8$ $0.8$ $0.46$ $1.027$
$10^{10}$ $1/100$ $1$ $1.5$ $0.50$ $0.6367$
$10^{3}$ $1/85$ $0.9$ $0.9$ $0.50$ $1.218$
$4\cdot 10^{3}$ $1/85$ $0.9$ $0.9$ $0.50$ $1.176$
We remark that the lower bound on $T$ in Theorem B.1 is slightly stronger than the bound we desire for Theorem 1.2. In fact one could weaken the bounds on $T$ in Theorem B.1 to get marginally better values for $K$. However, we chose to keep this lower bound unchanged as to maintain consistency between this paper and [4].